# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 2.4 Time Continuous Recurrent Neural Networks
To perform the forward pass of the network, the differential equation above is solved in the interval $t=t_{0}$ to $t=t_{1}$, subject to the initial condition of our input. In general, letting $\mathbf{I}(t)$ denote the input to the CTRNN at time $t$, we solve Eq. (2.1) in the interval $[t_{0},t_{1}]$ subject to the initial condition $x(t_{0})=\mathbf{I}(t_{0})$ for which $\sigma$ denotes a nonlinear activation function that is continuous and bounded (e.g., sigmoid). CTRNNs are more expressive than traditional networks due to the system of differential equations underlying their hidden state. These differential equations can be optimized through various numerical differential equation solvers, allowing CTRNNs to trade speed for accuracy in regions where it is necessary.
@@ 3 Theoretical Framework
[20] proposed a new framework for a quantum residual block in which they define a quantum residual operator applied to the entangled state |$\ket{\varphi_{0}}\otimes|0\rangle$. We utilize this approach to derive a quantum liquid network. A traditional residual block encoded into a quantum neural network is of the form:
|$\ket{\psi_{t+1}}=\ket{\psi_{t}}+\mathcal{U}\ket{\psi_{t}}.$ (3.1)
This approach does not retain the benefits of quantum computing. Instead, the authors of [20] present a quantum residual block that leverages the probabilistic nature of quantum computing to further enhance the capability of a residual block. Let $\mathcal{H}_{1}=\mathbb{C}^{2^{n}}$ and $\mathcal{H}_{0}=\text{span}\left\{\ket{0}\right\}$ denote Hilbert spaces. Consider an initial quantum state |$\ket{\varphi_{0}}\in\mathcal{H}_{1}$ and an auxiliary qubit |$\ket{0}\in\mathcal{H}_{0}$ as illustrated in Figure 3. The implementation of the quantum residual operator is realized in the subspace of the ancilla qubit. Once the residual operator is applied to the initial state, one can then obtain the results within the subspace $V=\left\{|x\rangle\langle x|\mid|x\rangle\in\mathcal{H}_{0}\right\}$ [20].
Inspired by [20], we develop an LQNet and a CTRQNet. First, we apply either a Hadamard gate to |$\ket{0}$ or a corresponding transformation to the state |$\ket{\varphi_{0}}$ depending on whether we are in the encoding stage or parameterized stage. We then apply a CNOT gate to the qubits using |$\ket{0}$ as the control qubit and |$\ket{\varphi_{0}}$ as the target qubit. From there, we apply another Hadamard to the quantum state whose output is measured in the auxiliary qubit subspace $\mathcal{H}_{0}$. The result of this operator is then added to the quantum state lying in the subspace of the first qubit. A final operation is performed on the state lying in $\mathcal{H}_{0}$ to realize the residual connection. Rewriting in a more succinct notation, we propose a Quantum Liquid Neural Network as follows :
Consider the entangled state |$\ket{\varphi_{0}}\otimes\ket{0}$. Then the output of the residual connection in the encoding stage(Eq. 3.1) can be rewritten in the following form:
@@
where $F:\mathcal{H}_{1}\rightarrow V$ denotes the action of the quantum residual block acting on classical data |$\ket{x}\in\mathcal{H}_{0}$ mapping to the subspace $V=\left\{|x\rangle\langle x|\mid\ket{x}\in\mathcal{H}_{0}\right\}$. We can think of the mapping $F$ as a sequence or composition of unitary operators
@@
where for the initial state |$\ket{\varphi_{0}}$, $U\triangleq\ket{\varphi_{0}}\bra{\varphi_{0}}$ |, CNOT denotes the controlled-not gate, $H$ denotes the Hadamard gate, and $I$ denotes the identity matrix. Note that $F$ is unitary by construction. Similarly, we define $\tilde{F}:\mathcal{H}_{1}\rightarrow V$ by
@@
where |$\ket{\psi_{t}}$ denotes the evolution of initial quantum state $|\varphi_{0}\rangle$ on layer $t$. ${\tilde{F}(\ket{\psi_{t}},\theta)}$ is equivalent to $W(\theta)\ket{\varphi_{0}}$, in $W(\theta)\ket{\varphi_{0}}\triangleq\mathcal{W}(\theta)$. With our definition of the residual block $\tilde{F}(\psi_{t},\theta)$, we have that in the limit, a model equation for quantum NODEs as follows
$\frac{d{\varphi}(t)}{dt}={\tilde{F}(\ket{\psi_{t}},\theta)},$ (3.6)
@@
$\frac{d{\varphi}(t)}{dt}=-\bigg{[}\tau^{-1}+{\tilde{F}(\ket{\psi_{t}},\theta)}\bigg{]}\varphi(t)+{\tilde{F}(\ket{\psi_{t}},\theta)}$ (3.7)
@@
${\varphi}(t+\Delta t)=\frac{{\varphi}(t)+\Delta t{\tilde{F}(\ket{\psi_{t}},\theta)}}{1+\Delta t\big{(}\tau^{-1}+{\tilde{F}(\ket{\psi_{t}},\theta)}\big{)}}$ (3.8)
@@
$\frac{d{\varphi}(t)}{dt}=-\tau^{-1}\varphi(t)+{\tilde{F}(\ket{\psi_{t}},\theta)},$ (3.9)
where the quantity ${\tilde{F}(\ket{\psi_{t}},\theta)}$ replaces $W(\sigma(\textbf{x}(t)))$ and is equivalent to the internal state of the output units of the network (cf. Theorem 2.3). This new equation (Eq. 3.9) forms the proposed model for CTRQNets.

### v2

@@ 2.4 Time Continuous Recurrent Neural Networks
To perform the forward pass of the network, the differential equation above is solved in the interval $t=t_{0}$ to $t=t_{1}$, subject to the initial condition of our input. In general, letting $\mathbf{I}(t)$ denote the input to the CTRNN at time $t$, we solve Eq. (2.1) in the interval $t_{0},t_{1}]$ subject to the initial condition $x(t_{0})=\mathbf{I}(t_{0})$ for which $\sigma$ denotes a nonlinear activation function that is continuous and bounded (e.g., sigmoid). CTRNNs are more expressive than traditional networks due to the system of differential equations underlying their hidden state. These differential equations can be optimized through various numerical differential equation solvers, allowing CTRNNs to trade speed for accuracy in regions where it is necessary.
@@ 3 Theoretical Framework
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]